if x-a/b + x-b/a = b/x-a + a/x-b slove for x............

Grade:12th pass

3 Answers

Vaas Volker
5 years ago
Given: x-a/b +x-b/a = b/x-a + a/x-b
a(x-a)+b(x-b)/ab = b(x-b) + a(x-a)/(x-a)(x-b)
Here Numerator on both sides are Equal so They Cancel Out, We Get
1/ab = 1/(x-a)(x-b)
(x-a)(x-b) = ab
x2-(a+b)x + ab = ab
x{x-(a+b)} = 0
Therefore X=0 or X-(a+b) = 0
X= 0 or (a+b)
mamta
5 years ago
there is (a-b) also in option …
 
Aditya Gupta
5 years ago
note that vaas has missed one solution which is:
x= (a^2+b^2)/(a+b).
so basically there are three solutions:
x= (a^2+b^2)/(a+b), 0, a+b